# Cross-modal Associations in Vision and Language Models: Revisiting the bouba-kiki effect

### v1

@@ 3 Methodology
Cross-modal associations are assessed through prompting two different versions of CLIP (Radford et al.,, 2021), specifically, we use a ResNet-50 and ViT version. We deem this as reasonable since state-of-the-art VLMs such as Molmo (Deitke et al.,, 2024), LLaVA (Liu et al.,, 2023), BLIP2 (Li et al.,, 2023), and InternVL (Chen et al.,, 2024) commonly use pre-trained frozen vision models and train lightweight alignment layers to align visual features with existing linguistic embeddings present in (large) language models (Liu et al.,, 2024). The ViT version of CLIP particularly often forms the basis for many current state-of-the-art VLMs (BLIP2, Molmo, InternVL, LLaVa, i.a.).  [^1] The source code and data used will be available upon publication.
@@
Models are probed through combining images and labels, the latter are pseudowords originating from four different sources. First, two sets of ’original’ labels are used that have been traditionally used the most in human studies (bouba-kiki and maluma-takete) such that we can explicitly compare human results to those of VLMs. Second, we borrow English adjectives from Alper and Averbuch-Elor, (2023) that are 10 synonyms of ’sharp’ and ’curved’. These serve as a baseline informing us whether the models can, in principle, make correct cross-modal associations. This differs from Alper and Averbuch-Elor, (2023), who use adjectives to create a vector to calculate a geometric score. Third, two-syllable labels were constructed following Nielsen and Rendall, (2013), using previously established cross-modal patterns in English. Sonorant consonants (M, N, L) and rounded vowels (OO, OH, AH) tend to match curved shapes, while plosive consonants (T, K, P) and non-rounded vowels (EE, AY, UH) align with jagged shapes. Combining these yields 36 syllables (e.g., loo, nah, kee, puh), categorised into four types: sonorant-rounded (S-R), plosive-rounded (P-R), sonorant-non-rounded (S-NR), and plosive-non-rounded (P-NR). These were paired into two-syllable pseudowords, loosely replicating the task in Nielsen and Rendall, (2013). For analysis, we focus on ’pure’ pseudowords that are either fully round-associated S-R-S-R or fully sharp-associated P-NR-P-NR, yielding 162 labels. Fourth, VLMs are tested on pseudowords generated using the method from Alper and Averbuch-Elor, (2023). Each pseudoword follows a three-syllable structure, where consonants (P, T, K, S, H, X, B, D, G, M, N, L) are combined with vowels (E, I, O, U, A) [^2] The letter A is included per Alper and Averbuch-Elor, (2023) despite that it is typically not regarded to evoke cross-modal associations in humans.. Pseudowords repeat the first syllable at the end (e.g., kitaki, bodubo). Only ’pure’ items—composed entirely of syllables from a single class—are used, excluding mixed forms like ‘kiduki’.
The linguistic inputs are comprised of a label of interest–which should induce certain preferences–and a prompt (‘The label for this image is <label>’) such that embedding the label is closer to the models’ natural objective. Provided that the pseudowords are generated anew, it is unlikely that they are known to the models. Ten different prompts are used (see Appendix A) to ensure that the results are robust and not an artefact of peculiarities in the prompts. These prompts are sourced from Verhoef et al., (2024), Alper and Averbuch-Elor, (2023), or are newly created such that half of the labels appear as a noun, and the other half as an adjective.
@@
The images fed to the models are either curved or jagged. Some of these images are sourced directly from previous works involving human participants (Köhler,, 1929, 1947; Maurer et al., 2006a, ; Westbury, 2005b, ), while others are specifically generated to test cross-modal associations in VLMs. These were inspired by the method described in Nielsen and Rendall, (2013) and already used by Verhoef et al., (2024). The generated images were created by randomly distributing points within a circle and then connecting them sequentially using curved lines for curved images and straight line segments for jagged images. Hence, this method generated image pairs that subtly differ only in features that are seemingly important for cross-modal effects. This is different from Alper and Averbuch-Elor, (2023), who generated images using Stable Diffusion, leaving less experimental control, which may distort which features cause the observed effects. An overview of all image pairs is displayed in Appendix B.
@@
All our analyses use Bayesian Regression Models as implemented in the brms package (Bürkner,, 2021) in R (R Core Team,, 2024). We fit models (4 chains of 4000 iterations and a warmup of 2000) to predict the proportion of correct guesses given a word_type with fixed effects of model, prompt, image pair, or label pair. The exact model formulas are displayed under each figure. Interpretation of our visualisation is straightforward: an effect is significant if the posterior means and their credible intervals are above chance level.
@@ 4 Probing through probabilities
To assess the preferences of the ViT and ResNet versions, we extract probabilities for all possible labels (i.e., syllables and pseudowords) conditioned on an image. For each image, we consider the label with the highest probability as the model’s preference (left in Figure 1). This is comparable to how Nielsen and Rendall, (2013) tested for human preferences. While using the probability data for all labels would have been possible, previous work demonstrated that using only the best-matching label yielded a small bouba-kiki effect, which was not the case when all probabilities were considered (Verhoef et al.,, 2024); as such, we use the most promising method. Importantly, we follow Ćwiek et al., (2022) who rightfully only treated human responses as bouba/kiki-congruent when participants matched bouba to a round shape, and crucially, also matched kiki to a spiky shape. In our work, this means that an image pair of a sharp and round shape must be matched with a congruent label.
Figure 2: The proportion of congruent responses for matching both images of an image pair correctly ($Match=1$). A result in which models consistently match images above chance (the grey dashed line) across prompts would suggest the presence of cross-modal associations. This is only the case for the English adjectives, which function as a baseline. Model: $Match\sim 1+Word\_type+(1+Word\_type|Prompt)$. Diamonds are descriptive means and the dots are posterior means.
@@ Results
First, we consider the English adjectives which serve as a baseline and do not inform about the bouba-kiki effect. Figure 2 reveals that these are successfully matched to images with round or sharp features. This result is consistent across prompts and is also somewhat expected. Yet, performance of roughly 80% across prompts and models could also be considered low given the clear-cut distinctions between images and adjectives. The primary labels of interest (i.e., the pseudowords) can not be robustly and correctly matched to images with corresponding features by either model version. Despite some variability across prompts and models, with some combinations leading to above-chance performance (e.g., ViT, Nielsen and Rendall,’s pseudowords, and This drawing is <label> or ResNet, Alper and Averbuch-Elor,’s pseudowords, and A picture of a <label> object), the descriptive means are mostly at chance level. Hence, we conclude that neither model displays clear cross-modal preferences. Interestingly, the individual modalities are in principle separable by both models (subsection A.1 and subsection B.1), suggesting that the difficulty lies in combining the information from multiple modalities. A look into which labels are mostly chosen reveals that model predictions mostly collapse onto a few frequently picked labels, regardless of the image input (Appendix C), strengthening our observation that no cross-modal associations influence predictions.
@@ 5 Beyond behavioural observations
In addition to using probability data, we are interested in knowing whether the predictions are wrong for the right reasons, i.e., do they not attend to curved or jagged regions when they see our images? As such, we further analyse the behaviour and preferences of models when associating an image with a label, using Grad-CAM, a technique from the model interpretability literature (Selvaraju et al.,, 2020). Grad-CAM provides a visual explanation of a model prediction by computing the gradients of the target class score, for which we use the cosine similarity between the label and image embeddings, with respect to the feature maps in the final convolutional layer or the final attention block. For this, we use public implementations of Grad-CAM (Zakka,, 2021; Mamooler,, 2021; Chefer et al.,, 2021). The values in these feature maps (typically displayed as a heatmap) can be interpreted as the importance or contribution of a specific region in the image to the prediction of a model (see Appendix D for visualisations of a random set of examples).
@@
This technique enables us to very closely mimic the cross-cultural study conducted by Ćwiek et al., (2022) in which human participants were shown the two classical bouba-kiki shapes and listened to the spoken words bouba or kiki. Hereafter, they selected which of the two shapes they thought corresponded to the word. To simulate this experiment with VLMs, we concatenate all image pairs into a single image containing a curved and jagged shape (n=17). We then identify for each label–which defines the expected target–which image region (left or right) receives the most attention from a model using Grad-CAM. While the generated heatmaps visually indicate the attention patterns, we use the sum of attention allocated to each curved and jagged part of the image to quantify the attention. Taking the total attention allocated to each part, instead of looking at very specific regions, aligns with how humans perceive faces, objects, and words holistically (Taubert et al.,, 2011; Zhao et al.,, 2016; Wong et al.,, 2011). We compare the sum of importance values in the expected region, given a pseudoword of a particular category, with the sum of importance of those in the non-expected region for each image and text label. Based on this comparison, we compute the percentage of trials in which the model consistently focuses more on the expected region than the non-expected region (right in Figure 1). The way images are concatenated are balanced such that the target appears eight or nine times on the left or right. Doing this eliminates a possible model preference for objects on the left or right. Initial analyses across models, prompts, labels, and images revealed that both models are relatively consistent in their predictions, regardless of the target location, with ResNet being consistent for $77.0\%$ of the predictions and the ViT variant $73.5\%$ (Appendix D).
Figure 3: The proportion of correctly matched responses for both labels of a label pair ($Correct=1$) given an image pair using Grad-CAM. The green line indicates human ‘performance’ reported in (Ćwiek et al.,, 2022). The grey line shows the chance level (25%) as the model must map ‘bouba’ and ‘kiki’ correct. Model: $Correct\sim 1+LabelPair+(1+LabelPair|Prompt)$. Diamonds are descriptive means and the dots are posterior means.
@@
Using Grad-CAM, we compare the model performances to that of English-speaking participants as reported by Ćwiek et al., (2022). While they only investigated the well-known label pair bouba-kiki, another pair, maluma-takete, combined with the two images shown in Figure 1, was first described by Wolfgang Köhler (Köhler,, 1929, 1947) and sparked wider interest in cross-modal associations. We only use these two label pairs here since the larger pseudoword set does not contain paired labels. However, our image generation process, which only varies how points are connected, allows evaluating across all image pairs, beyond just the original images. Specifically, we evaluate whether models attend more to the expected region (e.g., the sharp shape for "kiki" and "takete" or the round shape for "bouba" and "maluma"). A prediction is considered congruent only if both labels are correctly matched to their corresponding shapes. This setup closely mirrors human experiments and, in the case of the bouba-kiki label pair, enables direct comparisons.
@@
Extending the analyses beyond the original word pairs, we now assess all pseudowords to examine whether CLIP displays a fundamental association between syllables and shapes. The proportion of pseudoword-label responses displayed in Figure 4 reveals that there is again some variability across prompts, but model preferences are relatively consistent. Using this method, the ResNet variant displays a general preference to attend to round shapes across all word types (note that not attending to a jagged shape if the label is of the jagged category means the model attends to the round shape). Provided that this is also the case for the English adjectives, it is unsurprising that the pseudowords do not display a bouba-kiki effect. The ViT variant is relatively consistent and can reliably distinguish shapes for English adjectives. Yet, using experimental pseudowords collapses performance to chance levels, indicating the absence of human-like associations between shape features and syllables or characters. More evidence against the presence of cross-modal associations in CLIP comes from a qualitative inspection of several attention patterns. This reveals that attention patterns mostly do not focus on sharp edges or round attributes of images, but instead mostly focus on the centres of shapes or background areas (Appendix D). The latter is similar to earlier findings by Darcet et al., (2024) who identified that ViT networks create artefacts at low-informative background areas for computation purposes rather than describing visual information. Both observations are in contrast with what, at its core, is required for a bouba-kiki-like effect.
Figure 4: The proportion of correct matches (CorrectProportion) given a word type and category (i.e., curved or sharp). A cross-modal association is indicated when a model consistently matches images above chance (grey line) for both categories–observed only with the English adjectives used for comparison. Model: $CorrectProportion\sim 1+Word\_type+Category+(1+Word\_type+Category|Prompt)$. The diamonds are descriptive averages and the dots are posterior means.
@@ 6 Discussion and conclusion
This work investigated whether vision-and-language models, specifically two versions of CLIP (ResNet and ViT), exhibit human-like patterns of cross-modal association, as reflected in the bouba-kiki effect. We aimed to rigorously evaluate the presence or absence of this phenomenon in model behaviour, using methodological approaches closely aligned with human experiments and introducing the use of interpretability tools in this domain to probe internal representations. Contrary to prior work suggesting VLMs may encode bouba-kiki-like associations (Alper and Averbuch-Elor,, 2023), our findings reveal little to no evidence that CLIP models, under these experimental conditions, exhibit consistent human-like mappings between pseudowords and visual shape features. This lack of alignment suggests that, despite their impressive performance across many downstream vision-language tasks, models like CLIP do not appear to internally represent cross-modal associations in a cognitively grounded manner. The ResNet variant, which showed an overall preference for round shapes, did not demonstrate behaviour consistent enough to qualify as a bouba-kiki effect by human standards. The ViT-based CLIP, widely used as a foundational component in many current multimodal systems, showed even weaker alignment. Together, this raises questions about how cross-modal grounding is encoded or inherited in many larger state-of-the-art architectures.
@@
Our findings complicate earlier claims about the presence of sound-symbolic associations in model embeddings. While Alper and Averbuch-Elor, (2023) found strong evidence for a bouba-kiki effect using embedding similarity scores in CLIP and Stable Diffusion, Iida and Funakura, (2024) used the same method with Japanese VLMs, but failed to pick up robust bouba-kiki-like mappings, even though Japanese is rich in sound-symbolism (Dingemanse,, 2012). Instead, the embedding similarity scores seem to reflect language-specific regularities rather than general cross-modal preferences. Our approach deliberately avoided this by mirroring psycholinguistic testing paradigms with carefully controlled pairs of novel images. Corroborating other recent studies (Verhoef et al.,, 2024; Loakman et al.,, 2024), this empirically grounded probing method yielded opposite results, showing no evidence of a bouba-kiki effect. Whereas Verhoef et al., (2024) report limited, though not consistent, evidence for a bouba-kiki effect, our work differs since it uses a more comprehensive set of pseudowords, amending their analyses and suggesting that there is no general effect. We also ran the same tests using English adjectives, and in contrast to pseudowords, the models are able to make the expected mappings in that case. Furthermore, by employing Grad-CAM in the domain of cross-modal associations, we examined whether models visually attend to shape features that correlate with pseudoword form profiles. This enabled us to not only assess whether the models show human-like cross-modal associations, but also whether they do so for the right reasons by probing the decision-making pathways within the models. They largely did not attend to the expected shape features, which presents further evidence for the claim that VLMs may lack human-like cross-modal representations.
@@
The broader implication is that current VLMs, even those trained on massive paired datasets, lack a key component of human-like multisensory understanding: grounded, flexible, and intuitive mappings between sensory modalities. This is perhaps not surprising given the lack of embodied interaction in their training and reliance on statistical co-occurrence rather than perceptual salience. Given that the internal mechanisms of VLMs also inherently differ from humans, simply learning from aligned image-text pairs which are influenced by human-like preferences, including cross-modal associations, does not mean that they also learn the same preferences. Previous work on visual grounding (Jones et al.,, 2024), and shortcut learning, i.e., solving a test using unexpected non-human-like shortcuts (Schwartz and Stanovsky,, 2022; Mitchell and Krakauer,, 2023) argued similarly. The most straightforward difference is the issue of tokenisation. Unlike humans, who perceive words holistically (Wong et al.,, 2011), tokenisation fragments words, which distorts their representations and reduces potential human-like associations with visual shapes. The potential value of shared preferences between humans and machines, however, remains significant. Establishing alignment in human and machine understanding of (visual-auditory) form-meaning mappings and mutual understanding can enhance their interactions (Kouwenhoven et al.,, 2022), helping in creating AI systems that work effectively and resemble how humans process and communicate meaning. A potentially rich line of future work includes systematically exploring how design choices in VLMs affect the emergence of cognitively meaningful representations and consistently incorporating cross-linguistic studies to account for cultural variation (as demonstrated by Iida and Funakura, (2024)).

### v2

@@ 3 Methodology
Cross-modal associations are assessed by prompting two different versions of CLIP (Radford et al.,, 2021), specifically, a ResNet-50 and a ViT version. We deem this as reasonable since state-of-the-art VLMs such as Molmo (Deitke et al.,, 2024), LLaVA (Liu et al.,, 2023), BLIP2 (Li et al.,, 2023), and InternVL (Chen et al.,, 2024) commonly use pre-trained frozen vision models and train lightweight alignment layers to align visual features with existing linguistic embeddings present in (large) language models (Liu et al.,, 2024). The frozen ViT version of CLIP is particularly often the basis for many current state-of-the-art VLMs (BLIP2, Molmo, InternVL, LLaVa, i.a.). If CLIP, as a backbone, does not consistently exhibit human-like cross-modal associations, it is difficult to imagine how additional alignment layers can capture human-like representations without extensive fine-tuning. Moreover, CLIP enables us to extend existing approaches with an interpretability-based methodology that cannot be applied to proprietary models. [^1] The source data and code are available at https://osf.io/gqsv6/
@@
Models are probed through combining images and labels, the latter of which are pseudowords originating from four different sources. First, two sets of ‘original’ labels are used, which have been traditionally used the most in human studies (bouba-kiki and maluma-takete), allowing for explicit comparison of human results with those of VLMs. Second, we borrow English adjectives from Alper and Averbuch-Elor, (2023) that are 10 synonyms of ‘sharp’ and ‘curved’. These serve as a baseline informing us whether the models can, in principle, make correct cross-modal associations. This differs from Alper and Averbuch-Elor, (2023), who use adjectives to create a vector to calculate a geometric score. Third, two-syllable labels were constructed following Nielsen and Rendall, (2013), using previously established cross-modal patterns in English. Sonorant consonants (M, N, L) and rounded vowels (OO, OH, AH) tend to match curved shapes, while plosive consonants (T, K, P) and non-rounded vowels (EE, AY, UH) align with jagged shapes. Combining these yields 36 syllables (e.g., loo, nah, kee, puh), categorised into four types: sonorant-rounded (S-R), plosive-rounded (P-R), sonorant-non-rounded (S-NR), and plosive-non-rounded (P-NR). These were paired into two-syllable pseudowords, loosely replicating the task humans did in Nielsen and Rendall, (2013). For analysis, we focus on ‘pure’ pseudowords that are either fully round-associated (S-R-S-R) or fully sharp-associated (P-NR-P-NR), yielding 162 labels. Fourth, VLMs are tested on pseudowords generated using the method from Alper and Averbuch-Elor, (2023). Each pseudoword follows a three-syllable structure, where consonants (P, T, K, S, H, X, B, D, G, M, N, L) are combined with vowels (E, I, O, U, A). [^2] The letter A is included per Alper and Averbuch-Elor, (2023) despite that it is typically not regarded to evoke cross-modal associations in humans. Pseudowords repeat the first syllable at the end (e.g., ‘kitaki’, ‘bodubo’). Only ‘pure’ items—composed entirely of syllables from a single class—are used, excluding mixed forms like ‘kiduki’.
The linguistic inputs are comprised of a label of interest—which should induce certain preferences—and a prompt (‘The label for this image is <label>’) such that embedding the label happens at the sentence level and is closer to the models’ natural objective. Importantly, for each image, we only differ the pseudowords in question, so variation in probability must be a result of the pseudoword. Provided that the pseudowords are generated anew, it is unlikely that the models encountered them during training. Ten different prompts are used (see Appendix A) to ensure that the results are robust and not an artefact of peculiarities in the prompts. These prompts are sourced from Verhoef et al., (2024), Alper and Averbuch-Elor, (2023), or are newly created such that half of the labels appear as a noun, and the other half as an adjective.
@@
The images fed to the models are either curved or jagged. Some of these images are sourced directly from previous works involving human participants (Köhler,, 1929, 1947; Maurer et al., 2006a, ; Westbury, 2005b, ), while others are specifically generated to test cross-modal associations in VLMs. These were inspired by the method described in Nielsen and Rendall, (2013) and already used by Verhoef et al., (2024). The generated images were created by randomly distributing points within a circle and then connecting them sequentially using curved lines for curved images and straight line segments for jagged images. Hence, this method generated image pairs that subtly differ only in features that are seemingly important for cross-modal effects. This differs from Alper and Averbuch-Elor, (2023), who generated images using Stable Diffusion, leaving less experimental control and potentially distorting the understanding of which features cause the observed effects. An overview of all image pairs is displayed in Appendix B.
@@
All our analyses use Bayesian Regression Models as implemented in the brms package (Bürkner,, 2021) in R (R Core Team,, 2024). We fit models (using 4 chains of 4000 iterations and a warm-up of 2000) to predict the proportion of correct guesses given a Word_type with fixed effects for model, prompt, image pair, or label pair. The exact model formulas are displayed under each figure. The interpretation of our visualisation is straightforward: an effect is significant if the posterior means and their credible intervals are above the chance level.
@@ 4 Probing through probabilities
To assess the preferences of the ViT and ResNet versions, we extract probabilities for all possible labels (i.e., syllables and pseudowords) conditioned on an image. For each image, we consider the label with the highest probability to be the model’s preference (shown on the left in Figure 1). This is comparable to how Nielsen and Rendall, (2013) tested for human preferences. While using the probability data for all labels would have been possible, previous work demonstrated that using only the best-matching label yielded a small bouba-kiki effect, whereas using all probabilities did not (Verhoef et al.,, 2024); as such, we use the most promising method. Importantly, we follow Ćwiek et al., (2022), who rightfully treated human responses as bouba/kiki-congruent when participants matched bouba to a round shape, and crucially, also matched kiki to a spiky shape. In our work, this means that an image pair of a sharp and round shape must be matched with a congruent label.
Figure 2: The proportion of congruent responses for matching both images of an image pair correctly ($Match=1$). A result in which models consistently match images above chance (the grey dashed line) across prompts would suggest the presence of cross-modal associations. This is only the case for the English adjectives, which function as a baseline. Model: $Match\sim 1+Word\_type+(1+Word\_type|Prompt)$. Diamonds are descriptive means, and the dots are posterior means.
@@ Results
We begin by considering the English adjectives, which serve as a baseline and do not inform us about the bouba-kiki effect. Figure 2 reveals that these are successfully matched to images with round or sharp features. This result is consistent across prompts and is also somewhat expected. Yet, a performance of roughly 80% across prompts and models could also be considered low given the clear-cut distinctions between images and adjectives. The primary labels of interest (i.e., the pseudowords) can not be robustly and correctly matched to images with corresponding features by either model version. Despite some variability across prompts and models, with some combinations leading to above-chance performance (e.g., ViT, Nielsen and Rendall,’s pseudowords, and This drawing is <label> or ResNet, Alper and Averbuch-Elor,’s pseudowords, and A picture of a <label> object), the descriptive means are mostly at chance level. Hence, we conclude that neither model displays clear cross-modal preferences. Interestingly, the individual modalities are, in principle, separable by both models (A.1 and B.1), suggesting that the difficulty lies in combining information from multiple modalities. This is also visible when looking into which labels are most commonly chosen, revealing that model predictions do not differ much even though they are presented with different images (Appendix C). For the experimental pseudowords, models seem to rely on a few frequently selected labels (the percentage of unique labels for ResNet and ViT is $\approx 35\%$). This strengthens our observation that no robust syllable-level cross-modal associations influence predictions.
@@ 5 Beyond behavioural observations
In addition to using probability data, we are interested in knowing whether the predictions are wrong for the right reasons, i.e., do they fail to attend to curved or jagged regions when presented with our images? As such, we further analyse the behaviour and preferences of models when associating an image with a label, using Grad-CAM, a technique from the model interpretability literature (Selvaraju et al.,, 2020). Grad-CAM offers a visual explanation of a model’s prediction by calculating the gradients of the target class score. This involves using the cosine similarity between the label and image embeddings in relation to the feature maps from the final convolutional layer or the last attention block. In the case of ViT, we applied Grad-CAM to the last attention layer, which retains spatial information via its attention heads. We specifically focused on the attention from the [CLS] token to the image patches, similar to the approach in Caron et al., (2021) and following public implementations of Grad-CAM (Zakka,, 2021; Mamooler,, 2021; Chefer et al.,, 2021). Specifically, to identify which image regions contributed most to the decision, we computed the gradients of the target score with respect to the attention weights. These are averaged across heads, and we then extract the attention weights from the [CLS] token to the image patches by removing the [CLS] column. This is reshaped into a 2D spatial grid that acts as a feature map. The values in these feature maps (typically displayed as a heatmap) can be interpreted as the importance or contribution of a specific region in the image to the model’s prediction (see Appendix D for some visual examples).
@@
This technique enables us to very closely mimic the cross-cultural study conducted by Ćwiek et al., (2022) in which human participants were shown the two classical bouba-kiki shapes and listened to the spoken words bouba or kiki. Hereafter, they selected which of the two shapes they thought corresponded to the word. To simulate this experiment with VLMs, we concatenate all image pairs into a single image containing a curved and jagged shape (n=17). We then identify, for each label—which defines the expected target—which image region (left or right) receives the most attention from a model using Grad-CAM (see Figure 1). While the generated heatmaps visually indicate attention patterns, we quantify attention by summing the attention allocated to each curved and jagged part of the image. Taking the total attention allocated to each part, rather than focusing on very specific regions, aligns with how humans perceive faces, objects, and words holistically (Taubert et al.,, 2011; Zhao et al.,, 2016; Wong et al.,, 2011). Nevertheless, we also experimented with entropy and centroid-of-attention as quantifying measures, but none of these changed the results meaningfully (Appendix D. We compare the sum of importance values in the expected region for a given pseudoword within a particular category with the sum of importance values in the non-expected region for each image and text label. Based on this comparison, we compute the percentage of trials in which the model consistently focuses more on the expected region than on the non-expected region (as shown on the right in Figure 1). The way images are concatenated is balanced such that the target appears eight or nine times on the left or right. Doing this eliminates a potential model bias toward objects on the left or right. Initial analyses across models, prompts, labels, and images revealed that both models are relatively consistent in their predictions, regardless of the target location, with ResNet being consistent for $77.0\%$ of the predictions and the ViT variant $73.5\%$ (Appendix D).
@@
Using Grad-CAM, we compare the performance of both models to that of English-speaking participants as reported by Ćwiek et al., (2022). While they only investigated the well-known label pair bouba-kiki, another pair, maluma-takete, combined with the two images shown in Figure 1, was first described by Wolfgang Köhler (Köhler,, 1929, 1947) and sparked wider interest in cross-modal associations. We use only these two label pairs here, since the larger pseudoword set does not contain paired labels. However, our image generation process, which only varies how points are connected, allows evaluating across all image pairs, beyond just the original images. Specifically, we evaluate whether models attend more to the expected region (e.g., the sharp shape for ‘kiki’ and ‘takete’ or the round shape for ‘bouba’ and ‘maluma’). A prediction is considered congruent only if both labels are correctly matched to their corresponding shapes. This setup closely mirrors human experiments and, for the bouba-kiki label pair, enables direct comparisons.
@@
Figure 3: The proportion of correctly matched responses for both labels of a label pair ($Correct=1$) given an image pair using Grad-CAM. The green line indicates human ‘performance’ reported in (Ćwiek et al.,, 2022). The grey line shows the chance level (25%) as the model must map ‘bouba’ and ‘kiki’ correctly. Model: $Correct\sim 1+LabelPair+(1+LabelPair|Prompt)$. Diamonds are descriptive means, and the dots are posterior means.
Figure 4: The proportion of correct matches (CorrectProportion) given a word type and category (i.e., curved or sharp). A cross-modal association is indicated when a model consistently matches images above chance (grey line) for both categories—observed only with ViT and the English adjectives used for comparison. Model: $CorrectProportion\sim 1+Word\_type+Category+(1+Word\_type+Category|Prompt)$. The diamonds are descriptive averages, and the dots are posterior means.
@@
Extending the analyses beyond the original word pairs, we now assess all pseudowords to examine whether CLIP displays a fundamental association between syllables and shapes. The proportion of pseudoword-label responses displayed in Figure 4 reveals that there is again some variability across prompts, but model preferences are relatively consistent. Using this method, the ResNet variant displays a general preference to attend to round shapes across all word types (note that not attending to a jagged shape if the label is of the jagged category means the model attends to the round shape). Provided that this is also the case for the English adjectives, it is unsurprising that the pseudowords do not display a bouba-kiki effect. The ViT variant is relatively consistent and can reliably distinguish shapes for English adjectives. Yet, using experimental pseudowords collapses performance to chance levels, indicating the absence of human-like associations between shape features and syllables or characters. A qualitative inspection of several attention patterns provides additional evidence against the presence of cross-modal associations in CLIP. This reveals that attention patterns primarily do not focus on sharp edges or round attributes of images, but instead mainly focus on the centres of shapes or background areas (Appendix D). The latter is similar to earlier findings by Darcet et al., (2024) who found that ViT networks create artefacts at low-informative background areas for computational purposes rather than describing visual information. Both observations are in contrast with what, at its core, is required for a bouba-kiki-like effect.
@@ 6 Discussion and conclusion
This work investigated whether vision-and-language models, specifically two versions of CLIP (ResNet and ViT), exhibit human-like patterns of cross-modal association, as reflected in the bouba-kiki effect. We aimed to rigorously evaluate the presence or absence of this phenomenon in model behaviour, using two methodological approaches closely aligned with human experiments and introducing the use of interpretability tools in this domain to probe internal representations. Crucially, we posit that we can only speak of a consistent model preference when results across both methodologies are congruent with human-like preferences. Our first experiment, which used image-label probabilities to gauge model preference, showed no clear cross-modal preference for either model. The second experiment utilised Grad-CAM to simulate experimental conditions in human studies. Here, the ResNet variant, which showed an overall preference for round shapes, did not demonstrate behaviour consistent enough to qualify as a bouba-kiki effect by human standards. The ViT-based version of CLIP, widely used as a foundational component in many current multimodal systems, proved capable of matching shapes and English adjectives, but failed to do so for pseudowords, which have even weaker alignment than ResNet. Contrary to prior work suggesting VLMs may encode bouba-kiki-like associations (Alper and Averbuch-Elor,, 2023), our findings reveal little to no evidence that CLIP models, under these experimental conditions, exhibit consistent human-like mappings between pseudowords and visual shape features. This lack of alignment suggests that, despite their impressive performance across many downstream vision-language tasks, models like CLIP do not appear to represent cross-modal associations in a cognitively grounded manner internally. Together, this raises questions about how cross-modal grounding is encoded or inherited in many larger state-of-the-art architectures.
@@
Our findings complicate earlier claims about the presence of sound-symbolic associations in model embeddings. While Alper and Averbuch-Elor, (2023) found strong evidence for a bouba-kiki effect using embedding similarity scores in CLIP and Stable Diffusion, Iida and Funakura, (2024) used the same method with Japanese VLMs, but failed to pick up robust bouba-kiki-like mappings, even though Japanese is rich in sound-symbolism (Dingemanse,, 2012). Instead, the embedding similarity scores seem to reflect language-specific regularities rather than general cross-modal preferences. Our approach deliberately avoided this by mirroring psycholinguistic testing paradigms with carefully controlled pairs of novel images. Corroborating other recent studies (Verhoef et al.,, 2024; Loakman et al.,, 2024), this empirically grounded probing method yielded results opposite to those of other studies, showing no evidence of a bouba-kiki effect. Whereas Verhoef et al., (2024) report limited, though not consistent, evidence for a bouba-kiki effect, our work differs since it uses a more comprehensive set of pseudowords, amending their analyses and suggesting that there is no general effect. We also ran the same tests using English adjectives, and in contrast to pseudowords, the models are able to make the expected mappings in that case. Furthermore, by employing Grad-CAM in the domain of cross-modal associations, we examined whether models visually attend to shape features that correlate with pseudoword form profiles. This enabled us not only to assess whether the models exhibit human-like cross-modal associations, but also to determine whether they do so for the right reasons by probing the decision-making pathways within the models. They largely did not attend to the expected shape features, which presents further evidence for the claim that VLMs may lack human-like cross-modal representations.
@@
What causes this misalignment between humans and VLMs? The broader implication is that current VLMs, even those trained on massive paired datasets, lack a key component of human-like multisensory understanding: grounded, flexible, and intuitive mappings between sensory modalities. This is perhaps not surprising given the lack of embodied interaction in their training and reliance on statistical co-occurrence rather than perceptual salience. Still, it can be argued that the co-occurrences in the training data are full of human-like preferences, including cross-modal associations. So while we may expect VLMs to pick up on these associations, their internal mechanisms inherently differ from those of humans and do not necessarily learn the same preferences from aligned image-text pairs. Previous work on visual grounding (Jones et al.,, 2024), and shortcut learning, i.e., solving a test using unexpected non-human-like shortcuts (Schwartz and Stanovsky,, 2022; Mitchell and Krakauer,, 2023) argued similarly. However, interestingly, we observed that the individual modalities are, in principle, separable. This suggests that CLIP’s training objective (i.e., a sentence-level contrastive loss) and its method for aligning text and image modalities (i.e., cosine similarity) are the culprits. Neither seems to specifically promote learning relations between visual features and phonetic elements in language.
Another straightforward difference between humans and VLMs is tokenisation. Unlike humans, who perceive words holistically (Wong et al.,, 2011), tokenisation can distort word representations, reducing the potential for human-like associations with visual shapes. Inspection of the tokenised pseudowords, however, reveals that this does not occur frequently (‘Bouba’ becomes ‘bou’ and ‘ba’, ‘kiki’ is one token, ‘sepise’ becomes ‘sep’, ‘ise’, and ‘kaykuh’ becomes ‘kay’, ‘ku’, ‘h’. See section A.2 for more examples.) Most phonological structures are preserved that could match shape features. However, tokenisation may still split some pseudowords in ways that wouldn’t evoke the expected cross-modal associations in humans (e.g., ‘H’ in ‘OH’ might elicit jagged rather than curved associations). In the cases where tokenisation breaks pseudowords into syllables that also show no bouba-kiki effects in humans, the set of alternative pseudowords is large enough to contain options that could invoke these associations. If models had a genuine preference in the direction of a bouba-kiki effect, they would prefer those pseudowords that are tokenised as complete syllables, and retain the potential cross-modal association of interest, over those split into non-syllables in which the association may be corrupted. Our results show this is not the case. As such, the potential value of developing models with shared preferences between humans and machines remains significant. Establishing alignment in human and machine understanding of (visual-auditory) form-meaning mappings and mutual understanding can enhance their interactions (Kouwenhoven et al.,, 2022), helping in creating AI systems that resemble how humans process and communicate meaning. A potentially rich line of future work includes systematically exploring how design choices in VLMs affect the emergence of cognitively meaningful representations and consistently incorporating cross-linguistic studies to account for cultural variation (as demonstrated by Iida and Funakura, (2024)).